Ask your own question, for FREE!
Mathematics 8 Online
OpenStudy (anonymous):

A fair die is rolled 4 times. What is the probability that a 4 is obtained on at least one of the rolls? Round your answer to three decimal places.

OpenStudy (anonymous):

P(4 on at least one roll)=1-P(4 on no rolls) = 1-(5/6)^4 = 0.517747 = 52%

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
17 hours ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
20 hours ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Arizona: does anyone know the percentage rate of 30,000 and 40,000
1 week ago 0 Replies 0 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!